OpenStudy (anonymous):

Find the derivative f(t) = sqrt(t^2 +1), a = 3

OpenStudy (anonymous):

\[f'(t) = \frac{ 2t }{ 2\sqrt{t ^{2}+1} }\]

OpenStudy (anonymous):

had you understood how it become?

OpenStudy (anonymous):

I understand how you got 2t on the top, but on the bottom can you explain your steps? My algebra is apparently rusty, I didn't get that.

OpenStudy (anonymous):

it is like (2t^2 + 1)^1/2 when we diffrentiate number its power come in front and its power is become after subtracting 1 for example- derivative of x^3 is 3x^(3-1)

OpenStudy (anonymous):

Ohh I see you used the fast methods, my professor is making us use the definition derivatives

OpenStudy (anonymous):

So I multiplied it by \[\sqrt{(t+h)^2+1} + \sqrt{t^2 + 1} / \sqrt{(t+h)^2+1} + \sqrt{t^2 + 1}\] to remove the square root from the top, I see that after canceling a "h" I can get 2t on the top, but my math is wrong on the bottom and am not seeing how i can get what your got.

OpenStudy (anonymous):

i dont know about your method

OpenStudy (anonymous):